OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 4x + 3y = 2 3x + 2y = 1

OpenStudy (anonymous):

Write: I would use the elimination method, because neither x or y, has been isolated in terms of its counterpart.

OpenStudy (anonymous):

Now copy the tex I write

OpenStudy (anonymous):

I multiplied the second one by $$\frac{4}{3}$$, so that I got. $$4x+\frac{8}{3}y=\frac{4}{3}$$

OpenStudy (anonymous):

Then I subtracted the second one from the first one so that I got, $$\frac{y}{3}=\frac{2}{3}$$

OpenStudy (anonymous):

So I got that $$y=2$$

OpenStudy (anonymous):

Then I back substituted y, so that I got $$x=-1$$

OpenStudy (anonymous):

Now just copy everything, I wrote.

OpenStudy (anonymous):

ohh that's all ? thank you !

OpenStudy (zzr0ck3r):

if you don't like fractions, multiply the top by -3 and the bottom one by 4

OpenStudy (zzr0ck3r):

then no fractions!